OpenStudy (anonymous):

Input in standard form the equation of the given line. The line with m = -3/2 and b = 6

OpenStudy (mathmale):

good idea to start out by writing the general equation of a straight line in standard form. It is: Ax + By = C since you are given the slope and y-intercept of a line, your fastest route to an equation (any equation) for this line is the slope-intercept form, which is y=mx+b. Substitute the given values of m and b. Write out your result here.

OpenStudy (anonymous):

how would I set up the equation. 2=-3 + 6?

OpenStudy (mathmale):

PN: You are given m=-3/2 and b = 6. Just substitute those values into y=mx + b. Please give this another try.

OpenStudy (anonymous):

I dont know how to do that

OpenStudy (mathmale):

If y = m*x + b and m =-3/2 and b=6, then y = (-3/2)x + 6, or \[y=\frac{ -3 }{ 2 }x+6\]

OpenStudy (mathmale):

You will encounter problems like this one often in algebra and higher level math courses. any questions about this before we move on?

OpenStudy (anonymous):

-18/2?

OpenStudy (mathmale):

Remember, we're working with equations here. We're not trying to solve for x or y, but rather to write an equation of a line in standard form. If \[y=\frac{ -3 }{ 2 }x+6,\] then our next task is to re-write that so that it's in standard form: Ax+By=C.

OpenStudy (mathmale):

Try this: Multiply both sides of this equation by 2. (Multiply EVERY term by 2.) Please type out your resulting equation.